OpenStudy (anonymous):

Tamara is reviewing recent orders at her deli to determine which meats she should order. She found that of 1,000 orders, 450 customers ordered turkey, 375 customers ordered ham, and 250 customers ordered neither turkey nor ham. Based on these data, how many of the next 1,000 customers will order both turkey and ham? I think that you would take the turkey customers (450) added ham customers (375) and subtracted the neither customers (250) and solved for the probability, then use that for the next 1000 customers. But I'm not completely sure, and the book wasn't much help.

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):



OpenStudy (ranga):

450 + 375 - x + 250 = 1000 1075 - x = 1000 x = 1075 - 1000 x = 75 75 of the next 1000 customers will order both Turkey and Ham.

OpenStudy (anonymous):

I was almost right, thanks guys!

OpenStudy (ranga):

You are welcome.